# Elucidating the representation of images within an unconditional diffusion model denoiser

### v1

@@ Abstract
Generative diffusion models learn probability densities over diverse image datasets by estimating the score with a neural network trained to remove noise. Despite their remarkable success in generating high-quality images, the internal mechanisms of the underlying score networks are not well understood. Here, we examine a UNet trained for denoising on the ImageNet dataset, to better understand its internal representation and computation of the score. We show that the middle block of the UNet decomposes individual images into sparse subsets of active channels, and that the vector of spatial averages of these channels can provide a nonlinear representation of the underlying clean images. We develop a novel algorithm for stochastic reconstruction of images from this representation and demonstrate that it recovers a sample from a set of images defined by a target image representation. We then study the properties of the representation and demonstrate that Euclidean distances in the latent space correspond to distances between conditional densities induced by representations as well as semantic similarities in the image space. Applying a clustering algorithm in the representation space yields groups of images that share both fine details (e.g., specialized features, textured regions, small objects), as well as global structure, but are only partially aligned with object identities. Thus, we show for the first time that a network trained solely on denoising contains a rich and accessible sparse representation of images.
@@ 1 Introduction
Generative diffusion methods provide a powerful framework for sampling from probability densities learned from complex high-dimensional data such as images [1, 2, 3]. The key to their success lies in deep neural networks trained to estimate the score (the gradient of the log of the noisy image distribution), which is achieved by optimization on a denoising task. Training and sampling algorithms for these models have been extensively studied [4], but the score estimation properties that enable their spectacular image generation capabilities are not understood. Since denoising generally relies on distinguishing signal from noise, these score networks must somehow learn to identify and isolate patterns and structures found in their training data, so as to preserve them while eliminating noise. Here, we open the “black box” of a score network to study this internal representation.
@@
We examine a UNet [5] trained (unconditionally) for image denoising on the ImageNet dataset. Prior work [6, 7, 8, 9, 10] has shown that activations from such a model can be extracted to perform downstream tasks – such as classification and segmentation – with considerable success. We aim instead to understand what features arise in the internal space of the network. We find that the network computes a low dimensional sparse set of activations in the output channels of the middle (innermost) network block. This may be summarized with a vector comprised of the spatial averages of these channels, which provides a representation of the underlying clean image. We verify that this representation is stable when computed on an image contaminated with noise over a broad range of amplitudes, consistent with the denoising objective on which it is trained. We develop an algorithm for stochastic reconstruction, using a reverse diffusion algorithm conditioned on this representation vector. This decoding procedure recovers a sample from a set of images whose representation is the same as that of the target image. Visually, the reconstructed images are similar in terms of both local and global patterns, revealing both the commonality and diversity of attributes encoded in the representation.
@@
This representation exhibits a number of intriguing properties. Roughly, the channels summarized by the representation vector fall into two categories: non-selective channels, which capture common features present in many images; and selective channels, which are specialized for patterns that occur in only a small subset of images. As a consequence, the representation vectors of images lie within a union of low-dimensional subspaces, each of which is spanned by many of the common channels and a small subset of the specialized channels. We demonstrate that distances in the representation space are meaningful: The Euclidean distance between two representations predicts the symmetrized KL divergence between the two conditional distributions induced by the representations. Images whose representation vectors are similar are also semantically similar. Finally, we partition the images in the ImageNet validation set by applying a clustering algorithm to their representation vectors. The emergent clusters capture the general visual appearance of the images, sharing both fine details as well as global structure, but are only partially aligned with object category labels. Thus, the denoising objective alone, without any external conditioning, engenders learning of high level features that carry semantic information, albeit not necessarily object identities.
@@ 2 Implicit sparse image representation
The optimal solution is the conditional mean, which can be directly related to the score of the underlying noisy distributions using a relationship published in [11], but generally referred to as "Tweedie’s formula" [12, 13]:
@@
Thus, the trained network provides an approximation of the family of score functions for all $\sigma$. Reverse diffusion methods draw samples through iterative partial application of the learned denoiser, thereby using this approximate score to ascend the probability landscape [1, 2, 14, 3].
@@ 2.1 Spatial averages of activations
We adopt a convolutional UNet architecture [5], commonly used in diffusion models. The model consists of three main components: a set of encoder blocks (the downsampling path), a middle block, and a set of decoder blocks (the upsampling path) (see Fig 9 and Appendix A for more details). Our goal is to examine the activations of these hidden layers to understand how the noisy image is transformed to produce the score. To reduce the dimensionality, we examine the simplest summary of channel activations, $a_{j}(x_{\sigma})$, consisting of the spatial averages of all channels at the output layer of each block, notated by $\bar{a}_{j}(x_{\sigma})\in\mathbb{R}^{d_{j}}$. The components of the vector $\bar{a}_{j}$, which are always nonnegative due to half-wave rectification (ReLU) nonlinearities, carry information about the features encoded by the corresponding channels. Because of the convolutional nature of the network, the activations within each channel indicate the presence or absence of a feature in different spatial locations. Thus if a feature associated with channel $i\in(0,d_{j})$ is present somewhere in the input image, $\bar{a}_{j}(x_{\sigma})[i]$ will be positive, otherwise zero.
It is worth noting that the structure of the UNet architecture imposes certain properties on $\bar{a}_{j}$. Idealized convolutions are translation equivariant, hence the spatial averages of their activations are translation invariant w.r.t to the extracted features and do not carry information about location. For ergodic processes, these averages are approximations of statistical moments (expected value of nonlinear functions), and such measurements have been used for representation of visual texture [15, 16, 17, 18]. However, translation equivariance is imperfect for UNets (and most other convolutional networks) since it is violated by zero-padded boundary handling [19, 20]. This is especially prominent in the deeper blocks of the encoder, where the boundary influence encroaches on a larger spatial portion of the channels. In these cases, $\bar{a}_{j}$ can carry information about the location of features in addition to their presence. This effect coincides with the growth of the receptive field (RF) in the deeper blocks. As a result, $\bar{a}_{j}$ is expected to carry location information for larger features.
@@ 2.2 Denoising and channel sparsity
One of the difficulties in studying image representation in diffusion models is that it is not obvious where to look. We aim to locate the layers whose feature vectors, $\bar{a}$’s, well represent the clean image underlying the noisy input image. Such feature vectors would be rich enough to capture patterns and regularities of the image behind distortions caused by the noise in the input. Unlike Variational AutoEncoders (VAEs) [21], for which the encoder, bottleneck and decoder are well-defined and are assigned distinct roles through a dual objective function, the components of the UNet denoiser are only defined architecturally, and the optimization operates on a single end-to-end denoising objective. As a result it is not clear where the denoising takes place.
Some insight arises from considering the operation of denoisers designed in the pre-DNN era. Traditionally, image denoisers operate by transforming the noisy image to a latent space in which the true image is concentrated (sparse), and the noise is distributed (dense). Then the noise is suppressed and the image preserved. Finally, the latent representation is transformed back to the image space. This basic description captures the Wiener filter (operating in the frequency domain), and more modern thresholding denoisers which are typically applied within a multi-scale wavelet decomposition [22]. In all cases, more concentrated signal representations lead to better separation of noise and signal, hence superior denoising.
In the era of deep learning, neural networks optimized for denoising far outperform traditional solutions. Assuming the same principles apply, we expect to observe an increase in sparsity of the representation in portions of the network responsible for removing noise. To search for this locus, we measure the sparsity of $\bar{a}$ at the input and output layers of the blocks of a UNet trained on ImageNet. Here, sparsity of $\bar{a}\in\mathbb{R}^{d}$ is quantified by the normalized participation ratio (PR), the squared ratio $L_{1}$ and $L_{2}$ norms:
@@
Figure 1: Channel sparsity of input and output layers of the image denoising UNet. Inset histograms show normalized participation ratios (PR), of the spatially averaged input channels (orange) and output channels (green) for individual blocks. Encoder blocks exhibit increases in PR (i.e., decreased sparsity), while the middle block and decoder blocks exhibit increases in sparsity. Blocks $\{E_{1},D_{1}\}$ are not included since they have only one input/output channel, receptively. This is evidence that encoder blocks extract features to isolate noise and signal, and middle block and decoder blocks preserve those channels containing signal while suppressing those containing noise. (Notation: encoder blocks $\{E_{k}\}$, middle block ($M$), decoder blocks $\{D_{k}\}$, downsampling ($d$), upsampling ($u$),“skip” connections. See Appendix A)
Figure 1 shows distributions of sparsity for input/output layers of blocks in the UNet denoiser. Both encoding blocks offer a substantial decrease in sparsity of $\bar{a}$, but we observe a stark increase of sparsity in the middle and decoding blocks. This suggests that removal of noise, through suppression of channels whose activations are primarily carrying noise, starts at the middle blocks, and then continues in the decoder blocks. The decoder block $j$, hence, denoises the features present in $a_{j}$ passed from encoder $j$, conditioned on the denoised features coming from below. This is evidence that the encoder transforms the noisy image to a hierarchical representation in preparation for denoising, by extracting image structures learned from the training set. Thus, the representation lends itself to investigation at the output layers of the middle and decoder blocks, where the features are exposed after removal of noise.
@@ 2.3 Robustness of representation to noise
Figure 2: Stability of $\bar{a}$ across noise levels, for different network blocks. Plots show cosine similarity of $\bar{a}(x_{\sigma_{1}})$ and $\bar{a}(x_{\sigma_{2}})$, for $\sigma_{1}=0.5$, as a function of $\sigma_{2}$. $\bar{a}$ is most stable in the middle block (M). Note that $\bar{a}$ collapses as $\sigma$ falls to zero, for which the denoiser should compute the identity function.
To elucidate the relationship between $x$ and $\bar{a}$, we need first to clarify the effects of noise. Noise introduces two types of variability in $\bar{a}$: 1) $\bar{a}$ depends on the noise amplitude, $\sigma$, and 2) at any given noise level, $\bar{a}$ will also depend on the particular noise realization in $x_{\sigma}$. In the context of reverse-diffusion sampling algorithms, the former translates to the evolution of representation with time. Not surprisingly, the variance in $\bar{a}$ grows with noise level. To remove this variability, we take the mean of $\bar{a}$ across noise realizations, $\mathbb{E}_{z}\left[\bar{a}(x+\sigma z)\right]$. Fig 2 shows the effect of noise level on $\bar{a}$. Variability due to noise level depends on the block depth, but interestingly, $\bar{a}_{j}$ is most stable in the middle block: increased noise level mostly increases the amplitude but not the direction of $\bar{a}_{j}$. The noise resilience of $\bar{a}$ at the output layer of the middle block makes it a good candidate for study. Thus, for the remainder of this paper we focus on the representation in this layer, which we notate as:

### v2

# Unconditional CNN denoisers contain sparse semantic representations of images

@@ Abstract
Generative diffusion models learn probability densities over diverse image datasets by estimating the score with a neural network trained to remove noise. Despite their remarkable success in generating high-quality images, the internal mechanisms of the underlying score networks are not well understood. Here, we examine the image representation that arises from score estimation in a fully-convolutional unconditional UNet. We show that the middle block of the UNet decomposes individual images into sparse subsets of active channels, and that the vector of spatial averages of these channels can provide a nonlinear representation of the underlying clean images. Euclidean distances in this representation space are semantically meaningful, even though no conditioning information is provided during training. We develop a novel algorithm for stochastic reconstruction of images conditioned on this representation: The synthesis using the unconditional model is "self-guided" by the representation extracted from that very same model. For a given representation, the common patterns in the set of reconstructed samples reveal the features captured in the middle block of the UNet. Together, these results show, for the first time, that a measure of semantic similarity emerges, unsupervised, solely from the denoising objective.
@@ 1 Introduction
Generative diffusion methods provide a powerful framework for sampling from probability densities learned from complex high-dimensional data such as images (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020). The key to their success lies in deep neural networks trained to estimate the score (the gradient of the log of the noisy image distribution), which is achieved by optimization on a denoising task. Training and sampling algorithms for these models have been extensively studied (Croitoru et al., 2023), but the score estimation properties that enable their spectacular image generation capabilities are not understood. Since denoising generally relies on distinguishing signal from noise, these score networks must somehow learn to identify and isolate patterns and structures found in their training data, so as to preserve them while eliminating noise. Here, we open the “black box” of a score network to reveal this internal representation.
@@
Here, we examine a UNet (Ronneberger et al., 2015), trained unconditionally for image denoising. Prior work (Brempong et al., 2022; Yang and Wang, 2023; Xiang et al., 2023; Baranchuk et al., 2021; Mukhopadhyay et al., 2023) has shown that activations from such a model can be extracted to perform downstream tasks – such as classification and segmentation – with considerable success. We aim instead to understand and interpret those aspects of the internal representations of a denoiser that arise solely from the denoising task. We demonstrate that this network computes a low dimensional sparse set of activations in the output channels of the middle network block. This may be summarized with a vector comprised of the spatial averages of these channels, which provides a representation of the underlying clean image. We verify that this representation is stable when computed on an image contaminated with noise over a broad range of amplitudes.
@@
This representation exhibits a number of intriguing properties. Roughly, the channels summarized by the representation vector fall into two categories: non-selective channels, which capture common features present in many images; and selective channels, which are specialized for patterns that occur in only a small subset of images. As a consequence, for diverse datasets, the representation vectors lie within a union of low-dimensional subspaces, each of which is spanned by many of the common channels and a small subset of the specialized channels. We demonstrate that distances in this space are meaningful: Images whose representation vectors are similar are also semantically similar. This allows us to partition the images in the dataset by applying a clustering algorithm to their representation vectors. The emergent clusters capture the general visual appearance of the corresponding images, sharing both fine details as well as global structure, but are only partially aligned with object category labels.
Finally, we develop an algorithm for stochastic sampling, using a reverse diffusion algorithm conditioned on this representation vector computed from a target image. This procedure recovers a sample from a set of images whose representation is the same as that of the target image. Visually, these images are similar in terms of both local and global patterns, revealing both the commonality and diversity of attributes encoded in the representation. To quantify this, we show that the Euclidean distance between a pair of representation vectors strongly predicts the distance between a pair of conditional distributions induced by the representations. Thus, we show that the denoising objective alone, without any external conditioning, engenders learning of high level features that carry detailed semantic information.
@@ 2 Implicit sparse image representation
The optimal solution is the conditional mean, which can be directly related to the score of the underlying noisy distributions using a relationship published in Miyasawa (1961), but generally referred to as "Tweedie’s formula" (Robbins, 1956; Efron, 2011):
@@
Thus, the trained network provides an approximation of the family of score functions for all $\sigma$. Reverse diffusion methods draw samples through iterative partial application of the learned denoiser, thereby using this approximate score to ascend the probability landscape (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020; Kadkhodaie and Simoncelli, 2020).
@@ 2.1 Spatial averages of activations
We adopt a convolutional UNet architecture (Ronneberger et al., 2015), a smaller and more readily analyzed architecture than the most recent implementations, that nevertheless offers strong denoising performance. The model consists of three main components: a set of encoder blocks (the downsampling path), a middle block, and a set of decoder blocks (the upsampling path) (see Appendix A for details). We examine the activations of these hidden layers to understand how the noisy image is transformed to produce the score. To reduce the dimensionality, we consider the simplest summary of channel activations, $a_{j}(x_{\sigma})$, consisting of the spatial averages of all channels at the output layer of each block, notated by $\bar{a}_{j}(x_{\sigma})\in\mathbb{R}^{d_{j}}$. The components of the vector $\bar{a}_{j}$ are nonnegative (due to half-wave rectification (ReLU) nonlinearities), and carry information about the features encoded by their corresponding channels. Because of the convolutional nature of the network, the activations within each channel indicate the presence or absence of a feature in different spatial locations. Thus if a feature associated with channel $i\in(0,d_{j})$ is present somewhere in the input image, $\bar{a}_{j}(x_{\sigma})[i]$ will be positive, otherwise zero.
It is worth noting that the structure of the UNet architecture imposes certain properties on $\bar{a}_{j}$. Idealized convolutions are translation equivariant, hence the spatial averages of their activations are translation invariant w.r.t to the extracted features and do not carry information about location. For ergodic processes, these averages are approximations of statistical moments (expected value of nonlinear functions), and such measurements have been used for representation of visual texture Julesz (1962); Zhu et al. (1998); Portilla and Simoncelli (2000); Victor et al. (2017). However, translation equivariance is imperfect for UNets (and most other convolutional networks) since it is violated by zero-padded boundary handling, and downsampling operations. Both effects are more prominent in the deeper blocks of the encoder, which have undergone more downsampling, and for which the boundary influence encroaches on a larger spatial portion of the channels. In these cases, $\bar{a}_{j}$ can carry information about the location of features in addition to their presence. This effect coincides with the growth of the receptive field (RF) in the deeper blocks. As a result, $\bar{a}_{j}$ is expected to carry location information for larger features (Kadkhodaie et al., 2023; Kamb and Ganguli, 2024).
@@ 2.2 Denoising and channel sparsity
One of the difficulties in studying image representation in diffusion models is that it is not obvious where to look. We aim to locate the layers whose feature vectors, $\bar{a}$’s, well represent the clean image underlying the noisy input image. Such feature vectors would be rich enough to capture patterns and regularities of the image behind distortions caused by the noise in the input. Unlike Variational AutoEncoders (VAEs) (Kingma et al., 2013), for which the encoder, bottleneck and decoder are defined by their distinct assigned roles through a dual objective function, the components of the UNet denoiser are only defined architecturally, and the optimization is solely driven by a single end-to-end denoising objective. As a result it is not clear where the denoising occurs.
Some insight arises from considering the operation of denoisers designed in the pre-DNN era. Traditionally, image denoisers operate by transforming the noisy image to a latent space in which the true image is concentrated (sparse), and the noise is distributed (dense). Then the noise is suppressed and the image preserved. Finally, the latent representation is transformed back to the image space. This basic description captures the Wiener filter (operating in the frequency domain), and thresholding denoisers which are typically applied within a multi-scale wavelet decomposition (Milanfar, 2012). In all cases, more concentrated signal representations lead to better separation of noise and signal, hence superior denoising.
Figure 1: Channel sparsity of input and output layers of a UNet trained on ImageNet. Histograms show participation ratios (PR), of the spatially averaged input channels (orange) and output channels (green) for individual blocks. The middle block and decoder blocks exhibit increases in sparsity (i.e. reduction in PR). Blocks $\{E_{1},D_{1}\}$ are not included since they have only one input/output channel, receptively. This is evidence that encoder blocks extract features to isolate noise and signal, and middle block and decoder blocks preserve those channels containing signal while suppressing those containing noise. (Notation: encoder blocks $\{E_{k}\}$, middle block ($M$), decoder blocks $\{D_{k}\}$, downsampling ($d$), upsampling (uu),“skip” connections.) See Figure˜10 for other models.
In the era of deep learning, DNNs optimized for denoising far outperform traditional solutions. Assuming the same principles apply, we expect to observe an increase in representation sparsity in portions of the network responsible for removing noise. To search for this locus, we measure the sparsity of $\bar{a}$ at the input and output layers of each block of a UNet trained on ImageNet. Sparsity of $\bar{a}$ is quantified by the normalized participation ratio (PR), the squared ratio of $L_{1}$ and $L_{2}$ norms:
@@
Figure˜1 shows distributions of sparsity for input/output layers of blocks in the UNet denoiser trained on ImageNet64 dataset. Both encoding blocks offer a substantial decrease in sparsity of $\bar{a}$, but we observe a stark increase of sparsity in the middle and decoding blocks. This suggests that removal of noise, through suppression of channels whose activations are primarily carrying noise, starts at the middle blocks, and then continues in the decoder blocks. The decoder block $j$, hence, denoises the features present in $a_{j}$ passed from encoder $j$, conditioned on the denoised features coming from below. This is evidence that the encoder transforms the noisy image to a hierarchical representation in preparation for denoising, by extracting image structures learned from the training set. Thus, the representation lends itself to investigation at the output layers of the middle and decoder blocks, where image features are exposed after removal of noise. Figure˜10 shows this phenomenon in three additional UNet models trained on Texture, CelebA, and LSUN-Bedroom datasets.
Figure 2: Stability of $\bar{a}$ across noise levels, for different network blocks of a model trained on ImageNet64. Plots show cosine similarity of $\bar{a}(x_{\sigma_{1}})$ and $\bar{a}(x_{\sigma_{2}})$, for $\sigma_{1}=0.5$, as a function of $\sigma_{2}$. $\bar{a}$ is most stable in the middle block (M). Note that $\bar{a}$ collapses as $\sigma$ falls to zero, for which the denoiser should compute the identity function. See Figure˜11 for other models.
@@ 2.3 Robustness of representation to noise
To elucidate the relationship between $x$ and $\bar{a}$, we need first to clarify the effects of noise. The vector $\bar{a}$ depends on both the noise amplitude, $\sigma$, and the particular noise realization $x_{\sigma}$. In the context of diffusion sampling algorithms, the former translates to the evolution of representation with time. Not surprisingly, the variance in $\bar{a}$ grows with noise level. To remove this variability, we take the mean of $\bar{a}$ across noise realizations, $\mathbb{E}_{z}\left[\bar{a}(x+\sigma z)\right]$. Figures 2 and 11 show the effect of noise level on $\bar{a}$. Variability due to noise level depends on the block depth, but interestingly, $\bar{a}_{j}$ is most stable in the middle block: increased noise level mostly increases the amplitude but not the direction of $\bar{a}_{j}$. The noise resilience of $\bar{a}$ at the output layer of the middle block makes it a good candidate for study. Thus, for the remainder of this paper we focus on the representation in this layer, which we notate as:
@@
2.4 Denoising and channel selectivity
Figure 3: Channel selectivity. Left: Participation ratios for each channel over ImageNet. Distribution is bimodal, corresponding to channels that are highly specialized (and infrequently active) on left, and commonly used on right. Right: The panels show the set of images that maximally activate each of four specialized channels, revealing selectivity for rectangular periodic lattices (PR=0.19=0.19), a bird on a branch ($0.18$), cylindrical objects($0.19$), and dog faces ($0.26$). See Figure˜12 for other models.
In this section, we examine the components of the representation vector $\phi$, and their relationship to the content of the (clean) image $x$. For each channel, $i$, in the output layer of the middle block we quantify its (non-)selectivity using the participation ratio: $\frac{\|c(i)\|_{1}^{2}}{n\ \|c(i)\|_{2}^{2}}$, where $c(i)=\left(\phi({x_{1}}_{\sigma})[i],...,\phi({x_{n}}_{\sigma})[i]\right)$ is the concatenation of the $i^{th}$ entry of $\phi$ for all $n$ images in the dataset. This provides an estimate of the number of active channels, with smaller values indicating more selectivity. Figure 3 shows the distribution of selectivity for all $512$ channels in the middle block of a UNet trained on ImageNet64. The distribution is bimodal. Channels fall roughly into one of the two categories: selective channels that respond only to a small fraction of images in the set, and non-selective channels that respond to many images.
Specialized channels respond to specific features or patterns, and have negligible responses for most images in the dataset (Bau et al., 2020). Figure 3 shows sets of images that maximally excite four example specialized channels. The pattern shared across these images indicates the feature extracted by that channel. Since each image only contains a subset of patterns out of all the patterns present in the data distribution, only a handful of specialized channels are activated for each image. Figure 4 shows an example image along with the top three specialized channels it activates. Each channel is selective for a particular property of the image. Over the entire dataset, all channels are used, but different channels are used for different images.
More generally, channel selectivity predicts some statistical properties of the channels (Figure˜13). The marginal distribution of $\phi$ values in specialized channels is heavy-tailed, but common channels are closer to Gaussian. Specialized channels are spatially sparse while common channels are spatially denser. Additionally, PCA analysis on activation maps of specialized channels shows that they are highly concentrated in a few directions while the common channels are explained with more dimensions. Finally, less selective channels respond to a larger set of frequencies and orientations, and are the most stable w.r.t to noise level. These results imply that the common channels capture shared or common image features, such as brightness, global structure of a scene, etc.
2.5 Denoising and emergent semantic similarity
An immediate consequence of sparsity and channel selectivity is that the $\phi$’s lie in a union of subspaces in $\mathbb{R}^{d}$, whose dimensionality lies approximately in the range $[0.2d,0.3d]$ (from Figure 1). Each subspace corresponds to a combination of features that are likely to occur simultaneously (Figure 5). What do images whose $\phi$’s lie on the same subspace have in common? To examine this, we gathered images from the dataset with highest cosine similarity to a given target image in the representation space. Figures 5 and Figure˜14 show two sets of those images. Images that are closest in the $\phi$ are semantically similar, in stark contrast to images that are closest in the pixel space.
Emergent clusters Human-labeled object classes Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Representation vectors, $\phi$, are strongly concentrated within emergent clusters, less so with respect to object classes. Top Left: Euclidean distance between centroids of two clusters plotted against square root of their average variance along the line connecting the centroids, for 100 randomly selected cluster pairs. For all pairs, centroids are more than two standard deviations apart, and thus well-separated. t-SNE visualization of 7 randomly selected clusters. Top Right: Analogous plots for pairs of human-labeled classes, which are substantially more overlapping. Bottom: Each row contains images randomly selected from a cluster. Images within a cluster are visually similar and share global organization and semantic patterns, but they not necessarily from the same class. For example, dogs are clustered based on their pose, but not their breed. More examples in Appendix B.
The fact that pairwise Euclidean distances in the $\phi$ space are meaningful implies that similar images cluster together within the subspace. To test this, we applied a K-Means clustering algorithm (Lloyd, 1982) to the $\phi$ vectors computed from the ImageNet dataset (See Appendix B). Figure 6 shows that within clusters the $\phi$’s are well concentrated around centroids, and the centroids are well separated: centroids of adjacent clusters are at least 2 standard deviations apart. The $\phi$’s of human labeled classes, by comparison, are less well separated. As a result, we expect images from different classes to appear within the same cluster. Examples are shown in Figure 6.
Images within same clusters are semantically similar: they clearly share location-specific global scene structure, as well as location-nonspecific detailed elements (e.g., objects and their constituent parts, texture). However, in many cases, the similarity is not the object identity: a variety of circular objects appear in one cluster, dogs on a grass background in another, and a variety of bugs on a flat background in another. This reveals that the type of patterns that are useful for computing the score are not object identities. What connects images within the same cluster is "the gist of the scene" (Potter and Levy, 1969; Oliva and Torralba, 2001; Oliva, 2005; Oliva and Torralba, 2006; Sanocki et al., 2023), which is only partially consistent with human labeled class memberships. Remarkably, these emergent clusters arise solely from learning the score, by minimizing a denoising objective. The network is trained without labels, augmentations, or regularization to induce any specific type of similarity or grouping (e.g membership in the same class). Therefore proximity in the representation space defines a fully unsupervised form of similarity in the image space.